I don't know if this is the right section for this question:
Each week copies of business magazines are delivered to three different stores that have ordered 75 copies, 100 copies and 150 copies, respectively. How many copes should be packaged together so that no package should be opened during delivery?

Answer by checkley75(3666)

75,100,150 have a common factor which is 25
3*25=75
4*25=100
6*25=150
thus packages of 25 will do the job.